A cobb douglas production function for a new company is given by F(x,y)=50x3/5y2/5. Where x represents the units of labor and y represents the units of capital. Suppose the units of labor and the capital cost is $200 and $100 each respectively and the budget constraint is $30,000. Find the maximum production level for this manufacturer.


1
Expert's answer
2021-07-09T11:16:27-0400
"200x+100y=30000=>y=300-2x, x, y\geq 0"

Then


"F=F(x)=50x^{3/5}(300-2x)^{2/5}, 0\leq x\leq 150"

Find the first derivative with respect to "x"


"F'(x)=(50x^{3\/5}(300-2x)^{2\/5})'"

"=50(\dfrac{3}{5}x^{-2/5}(300-2x)^{2/5}-\dfrac{2}{5}(2)x^{3/5}(300-2x)^{-3/5})"

"=10x^{-2\/5}(300-2x)^{-3\/5}(3(300-2x)-4x)"

"=100x^{-2\/5}(300-2x)^{-3\/5}(90-x)"


"F'(x)=0=>100x^{-2\/5}(300-2x)^{-3\/5}(90-x)=0"

"=>x=90"

Critical numbers: "0, 90, 150"

"F(0)=0"


"F(150)=0"



"F(90)=50(90)^{3\/5}(300-90)^{2\/5}="

"=1500(1323)^{1/5}\approx6315.43"

The maximum production level is $6315.43 and occurs when x=90 units of labor and y=120 units of capital.